# Challenges and Requirements for Benchmarking Time Series Foundation Models

### v2

@@ 3.2 Indirect Information Leakage: Temporal Overlap of Correlated Train and Test Series
This temporal overlap of correlated train-test samples introduces a problematic dependency between the training and test data, violating the assumption of independence that is crucial for robust evaluation. As a result, the model gains access to information during training that would be inaccessible in a real-world scenario at the time of prediction, leading to inflated performance metrics. Therefore the explicit time period of a pattern and its actually encoded information gain importance for globally trained models such as TSFMs.
This issue is not merely theoretical, but has been demonstrated empirically. Rodrigo and Ortiz (2024) conducted an experiment isolating the influence of temporal overlap information leakage. They build a setup with highly correlated time series, such as metro, bus, road, and train usage in Madrid for the year 2024. They trained two models in a global manner with all series. One model was trained with a common train-test split date for all series. For the other model only the metro time-series was cutted at the train-test split date, keeping the other series fully in the train set. After training, they only predicted the test period of the metro series, which was never part of the training dataset. The second model with full access to the other correlated series during training achieved a Mean Absolute Error (MAE) of 248,24 compared to 439,14 to the first model with a common train-test split date across all series. This corresponds to a performance improvement of approximately 43.47% attributable solely to this indirect form of leakage, confirming that access to correlated and temporally overlapping external information significantly distorts evaluation results even without direct access to the target series.
@@
While the study by Rodrigo and Ortiz (2024) used time series where the complete training data is highly correlated and temporally overlapping with the complete test data, we replicate an TSFM training process in a more realistic scenario. We incorporate diverse data where only a small percentage of observations is affected by indirect leakage. To demonstrate the impact of temporal train-test leakage on model performance under these conditions, we conducted an experiment using a trained-from-scratch Time-Series Transformer (TST) (Zerveas et al., 2020). As we have seen in Section 3.1, any other pre-trained TSFM could be already affected by this form of information leakage. The encoder-only model shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) but is significantly smaller, enabling faster training saturation. We trained the model using 10 different random seeds for two different training sets [^9] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section C: one with a temporal overlap to the test set during the COVID-19 pandemic and one without such overlap. The first training set comprised the NN5 Daily and Tourism Monthly (Monash collection), and the American (Dow Jones), Brazilian (Bovespa), British (FTSE 100), Canadian (TSX), French (CAC 40), German (DAX) and Japanese (Nikkei 225) stock indices from 2015 to the end of 2019. The only difference in the second training set was that the stock index data spanned from 2016 to the end of 2020, comprising observations affected by the COVID-19 outbreak. No data points were shared directly between the training and test sets.
Table 2: Per-seed S&P 500 MAE breakdown for indirect leak from other countries stock indices vs no-leak experiments. The percentage difference shows how much better the model with leakage data performs than the non-leakage model. For each seed, the better MAE in displayed in bold.
Seed MAE Before Split MAE After Split
No Leak Indirect Leak % Diff No Leak Indirect Leak % Diff
410 230.53 227.72 1.22% 157.85 199.33 -26.28%
1425 219.68 178.09 18.93% 120.77 82.53 31.67%
1680 193.52 187.19 3.27% 80.74 81.22 -0.60%
1825 247.76 156.12 36.99% 110.47 70.36 36.31%
2287 231.19 206.28 10.78% 102.42 100.54 1.84%
3658 219.23 156.91 28.43% 117.25 84.62 27.84%
4013 224.75 193.16 14.06% 114.19 97.98 14.20%
4507 206.45 213.70 -3.51% 93.10 84.01 9.77%
8936 189.40 199.38 -5.27% 103.59 153.67 -48.35%
9675 206.30 182.99 11.30% 82.35 132.44 -60.82%
Mean 216.88 190.15 11.62% 108.27 108.67 -1.44%
Std Dev 17.15 21.88 21.10 38.83
We selected stock prices as our target variable as they are specifically hard to model appropriately in a long-term valid manner Goyal et al. (2024), suggesting that any performance gaps must come from external leaked information. For both models, we used the S&P 500 index as a zero-shot test set, which was not part of both datasets. Additionally, we chose a time-frame where a common cause had influence on many different time series and chose therefore the COVID-19 outbreak. In Figure 3, time series 1 shows the DAX index and time series 2 represents the S&P 500 index. Note how the COVID-19 outbreak is visible in both series.
Figure 4 displays the mean and standard deviation of the predictions for both models across the ten runs, and Table 2 shows the specific values for each seed. For almost all seeds, the model trained on data with temporal overlap to the test set predicted the initial market crash following the outbreak significantly better than the non-leakage model. In total, the temporal overlap model performed better eight out of ten times. Up to the first of May, the leakage model achieved an average MAE of about 190 compared to about 217 for the non-leakage model. For the remaining part, both models had a very similar average MAE, with the non-leakage model being slightly better (108.3 compared to 108.7 for the non-leakage model. When looking at the specific runs, shown in Table 2, the performance difference between the two models is up to about 37% in favor of the leakage model during the shock. Despite the temporal overlap representing only a small fraction of the overall training data, specifically seven time series out of 484 (or 1,600 out of 205,859 observations in total) (Table 6), this performance gap underscores the significant impact the small temporal overlap of correlated series can have on the predictive performance of a model.
4 TSFM Benchmark Requirements
@@
However, the paradigm shift behind TSFMs introduces new possibilities for information leakage. Unlike local models, TSFMs learn patterns globally. Consequently, strict temporal ordering within a single time series is no longer sufficient to guarantee a fair evaluation. To rigorously assess these models, we must address leakage across two distinct dimensions: the sample dimension and the temporal dimension (see Figure 5).
@@
Flexibility in Pre-Training Data (Training Dimension) The current lack of transparency surrounding TSFM pre-training corpora creates a substantial risk of information leakage as shown in Section 3.1. This growing confusion is evident in newer benchmarks and models that report results contaminated by data leakage—whether unintentional or intentional Aksu et al. (2024); Auer et al. (2025); Li et al. (2024b). Attempting to mitigate this risk by restricting pre-training data is both counterproductive and ineffective: such restrictions artificially limit the potential of foundation models and, given the opacity of large-scale pre-training corpora, our lineage analysis shows that it is almost impossible to rule out train-test sample overlaps. We therefore argue for full flexibility in pre-training data. A robust evaluation framework should be agnostic to pre-training data choices, allowing the use of any existing data to fully exploit the predictive capabilities of TSFMs. This perspective shifts the responsibility for preventing leakage from the TSFM developer to the benchmark designer, thereby necessitating evaluation datasets that are genuinely novel and capable of assessing generalization rather than memorization.
Strict Global Post-Training Test Periods (Inference Dimension) Indirect temporal dependencies between time series, originating from direct or indirect causal relationships, create another more subtle risk for information leakage. Preventing this risk by enforcing global cut-off dates for pre-training data to separate training from test data is not a viable solution to this challenge. First, it would require all researchers to tailor their training protocol to exactly this date and second, it prevents new time series patterns of current time periods after the cut-off from being trained into the models. We therefore argue for strict global post-training test periods. These test periods need to be flexibly determined based on the global last pre-training time point of all evaluated models, instead of forcing a pre-training cut-off. For a robust evaluation framework a strict temporal barrier must be enforced across the entire universe of data: if the last pre-training timestamp across all evaluated models is $t$, only observations from $t+1$ onwards can be present in the test set of a benchmark, ensuring that TSFMs rely solely on past dynamics.
Addressing these two dimensions (unrestricted training data and strict temporal inference separation) is non-trivial and often overlooked in current evaluation standards. We ask the community to develop ideas for new benchmark methodologies which are grounded in these two requirements. Possible directions could be constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data, just to name a few.
@@ 5 Conclusion
Remarkably, TSFMs appear to follow a development trajectory similar to that of LLMs in terms of both data and model size. Scaling laws seem to apply to TSFMs, and there are notable similarities in the evaluation challenges they face, particularly concerning potential information leakage. We highlight key evaluation challenges that arise from the emergence of TSFMs. By investigating the data lineage of 22 prominent TSFMs and reflecting on common TSFM training and evaluation strategies, we identified two significant sources of information leakage: direct information leakage by train-test sample overlaps through multi-purpose use of datasets and indirect information leakage by temporal overlap of correlated train and test series.
Current benchmark approaches from time series forecasting or machine learning do not mitigate all information leakage risks. Based on our investigations and existing methodologies, we propose two key requirements for information leakage free evaluations in the age of TSFMs: flexibility in pre-training data and strict global post-training test periods. Beyond the development of new benchmark methodologies, we identify several directions for future research . To date, the effect size of test-set contamination—when a test dataset is present in the pre-training corpus—has only been shown to be significant in isolated cases. Likewise, the magnitude of temporal leakage arising from correlated and overlapping time series remains largely unknown in real TSFM settings and is likely to vary across domains, types of global events, and other contextual factors. Carefully designed TSFM training regimes will be required to rigorously quantify these effects.
@@ 6 Data Availability
The resulting table of the TSFM data-lineage analysis is supplied in the Supplementary Information Section D. For the global pattern experiment in Section 3.2, the used data is available through the source code via zenodo: https://doi.org/10.5281/zenodo.18635478
@@ 7 Code Availability
The code to reproduce the experiment in Section 3.2 is provided via zenodo: https://doi.org/10.5281/zenodo.18635478
Appendix A Time Series Foundation Model Analysis
A.1 Multi-Purpose Use of Train and Test Data
This section showcases the varying training and testing dataset strategies of selected TSFMs. Table 3 reveals that there is virtually no standardized approach among TSFMs, with each model employing different combinations of datasets for train/test (or in-domain) datasets. While some datasets like Electricity (H) and Traffic Hourly (H) appear across multiple models, the coverage remains fragmented. Table 4 further illustrates this inconsistency in zero-shot evaluation practices. Even among datasets with five or more overlaps, the specific models testing on each dataset vary considerably. For instance, ETT2 (15T) is evaluated by Chronos, Moirai, and Lag-Llama, while Electricity (H) appears in Moirai, Moirai-MoE, and LLMTime.
Another notable case for the multi-purpose use of datasets is the subset Australian Electricity Demand from the Monash dataset collection (Godahewa et al., 2021), which is used for pre-training (e.g., Lag-Llama, Timer), train/test evaluation (e.g., Moirai), and zero-shot forecasting (e.g., Chronos, TimesFM). This multi-purpose use makes the dataset practically useless for benchmarking between models (e.g., it cannot be used to compare Chronos to Lag-Llama). To understand the problem of overlaps better, we provide Figure 6.
Also an interesting case is the Informer collection (Zhou et al., 2021), which is primarily used for zero-shot evaluations. However, some individual time series of the Informer collection, such as ETTh1, ETTh2, and ETTm1, have also been used for pre-training (e.g., in LagLlama) or train/test evaluation (e.g., in UniTime). This means that LagLlama or UniTime cannot be benchmarked against any other model using the Informer collection.
Taken together with investigations of Roque et al. (2025) that the selection of evaluation datasets can highly influence the final evaluation results, this strengthens the need for a more uniform benchmark approach.
Figure 6: Test set contaminations when comparing TSFM A with TSFM B with overlapping pre-training and test sets on the test set of TSFM A.
Table 3: Comparison of train/test evaluations of selected different models ordered by the total number of overlaps. For full results see Section D.
Dataset Chronos Moirai MOMENT Time-LMM
Electricity (H) T/T - T/T T/T
Traffic Hourly (H) - T/T T/T T/T
ETT1 (H) - - T/T T/T
Weather (Australia) (D) - T/T - T/T
M4 Monthly (M) T/T T/T - -
M4 Hourly (H) T/T T/T - -
M4 Daily (D) T/T T/T - -
M4 Weekly (W-SUN) T/T T/T - -
ETT1 (15T) - - T/T T/T
ETT2 (15T) - - T/T T/T
ETT2 (H) - - T/T T/T
CDC Fluview ILINet (W) - - T/T T/T
Table 4: Comparison of zero-shot evaluations of selected different models ordered by the total number of overlaps. Only datasets with 5 or more overlaps are shown. Models which have not included any of the shown datasets into their train/test evaluation are not displayed. For full results see Section D.
Dataset Chronos Moirai Moirai-MoE Lag-Llama LLMTime
Jena Weather (10T) - ZS ZS - ZS
ETT2 (15T) ZS ZS - ZS -
Electricity (H) - ZS ZS - ZS
ETT2 (H) ZS ZS - - -
ETT1 (H) ZS ZS - - -
ETT1 (15T) ZS ZS - - -
Traffic Hourly (H) ZS - - - ZS
Appendix B Leakage Investigations
B.1 Effects of Overlaps in Train-Test Samples
@@
This section collects different examples where the train-test sample overlaps unintentionally led to information leakage in recent benchmark papers. In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used in the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b). Based on our analysis of the benchmark results, this leads to an advantage of 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets. The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%. The performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section B.2). In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset which is included in every tested TSFM pretraining data Saravanan et al. (2024). Also in the energy domain, a paper benchmarked TSFMs on the Spanish dataset (among others), which is also included in the pretraining data of the tested TSFMs. Chronos achieved a sMAPE of 4.854 where the best non-TSFM model TiDE achieved a sMAPE of 8.102 Montet et al. (2025). This highlights the increasing difficulties for benchmarking TSFMs.
B.2 Intended Information-Leakage Experiment for Moirai Model
@@
Table 5: Information leakage in TSFM comparing Moirai with (MoiLeak) and without (Moi) information leakage during training phase for different model sizes. Metric is MAPE. Based on experiments by Aksu et al. (2024).
@@
The empirical analysis using the Moirai TSFM by Aksu et al. (2024) demonstrates the effects of information leakage between training and test sets on the predictive performance of TSFMs. The authors prepared a new pre-training and held-out evaluation dataset without any overlap and trained the Moirai architecture on it. A different Moirai model from a previous publication was taken which pre-training dataset contained 0.1% data of the newly defined held-out evaluation dataset leading to a deliberate small portion of information-leakage. The investigated data includes nine datasets tested on three different forecast horizons (short, medium, long). Not for all horizons data was available, so we filtered the results on 4 datasets where the results are available for all horizons to reasonably compare the metrics and model sizes, as shown in Figure 7. On short horizons, both models, leakage-free and leakage model, perform better with increasing model size while Moirai Leakage has on average 7.6% points lower MAPE score. At medium horizons, the leakage-free MoiraiL generally underperforms compared to other model sizes, exhibiting an average performance deficit of 32% points compared to the Moirai LeakageL model. On long horizons, all models perform significantly worse. However, Moirai Leakage achieves an average MAPE score that is 29% points lower than the leakage-free Moirai. Across all horizons, the biggest difference is between MoiraiL and Moirai LeakageL.
These results vividly demonstrate the significant impact that train-test sample overlap has on the performance of TSFMs. Larger model sizes especially seem prone to this.
@@
B.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
It may be theoretically possible that existing model evaluation results have already been influenced by overlaps of correlated train and test series. We analyse the overlapping time frames in the Monash Dataset  [^10] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 8. Test datasets shouldn’t lie in the same time period as other training datasets. As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021). The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if their are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis. Similarly, the Oikolab Weather dataset ranges from 2010 January to 2021 May, which is weather data located in Melbourne. The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020. This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example. Both datasets are included in the datasets of Moirai, Time-Moe, Sundial and Toto.

### v3

@@ 3.2 When Temporal Overlap Turns Correlation Into Leakage
The evaluation implication follows directly. A key strength of TSFMs lies in their ability to learn subtle patterns in a data-driven manner and transfer them to previously unseen series. At large scale, however, this same capability becomes a liability for evaluation. When correlated training and test series overlap temporally, a sufficiently flexible TSFM can transfer learned statistical structure from the former to the latter. Crucially, this does not reflect genuine forecasting skill: the model is not learning the generative dynamics of the test series, but exploiting correlation produced by a shared external driver. As illustrated in Figure 3, a TSFM trained on the DAX index through the COVID-19 crash absorbs the temporal signature of that event. When subsequently asked to forecast the S&P 500 over the same window, it can reproduce that shape — not by understanding US market dynamics, but because both indices were driven by the same global shock and are therefore highly correlated during that period. The key distinction from classical benchmarking is therefore not only which series appeared in training, but when those series were observed: contemporaneous observation of correlated series embeds information about the test period even when no test data point was ever directly seen. This temporal overlap violates the independence assumption fundamental to unbiased evaluation, inflating performance metrics in ways that do not reflect real-world forecasting ability.
This concern is not merely theoretical. Rodrigo and Ortiz (2024) provided a clean empirical demonstration using highly correlated public transport series from Madrid: metro, bus, road, and train ridership throughout 2024. The metro series was withheld entirely from training and served as the zero-shot forecasting target. For all remaining series, two training configurations were compared: in the first, a common train-test split date was applied, so that none of the correlated series extended into the metro test period; in the second, all non-metro series were kept in training up to the end of the evaluation window, overlapping temporally with the metro test period. Both models were then evaluated exclusively on the metro test period. The model whose correlated training series overlapped with the test period achieved a Mean Absolute Error (MAE) of 248.24 compared to 439.14 for the properly split model, a performance improvement of approximately 43% attributable solely to this indirect form of leakage. No metro data point was ever seen during training; the advantage came entirely from the shared temporal dynamics of the correlated series.
The Madrid experiment used a near-ideal setting for leakage: virtually all training series were highly correlated with the test target. The more pressing question for TSFM evaluation is whether temporal leakage remains detectable when the leaked signal is diluted among a large, heterogeneous training corpus. A controlled replication around the COVID-19 stock market crash of early 2020 suggests the answer is yes. To isolate the effect cleanly, a Time-Series Transformer (TST) (Zerveas et al., 2020) was trained from scratch, rather than relying on an existing TSFM that could already carry this form of leakage. The architecture shares multi-variate characteristics with recent models such as MOMENT or Chronos-2 (Goswami et al., 2024; Ansari et al., 2025) while remaining small enough for fast training saturation. Two almost identical training sets were constructed [^7] Details on the general setup, used datasets, scaling, hyperparameter-tuning and results can be found in Supplementary Information Section B.: both contained the same non-financial series (NN5 Daily and Tourism Monthly from the Monash collection) plus seven major stock indices (Bovespa, CAC40, DAX, DowJones, FTSE 100, Nikkei 225, TSX) [^8] All stock index data was retrieved from the website investing.com, differing only in the temporal range of the stock data. The first set cut the stock series at December 2019, before the pandemic; the second extended them through December 2020, encompassing the crash and partial recovery. In both cases, the S&P 500 was withheld entirely as the zero-shot test target. Stock prices were chosen deliberately: they are notoriously hard to model beyond short horizons (Goyal et al., 2024), ensuring that any systematic performance gap reflects leaked external information rather than improved modeling of intrinsic dynamics.
@@
Figure 4 displays the mean and standard deviation of the predictions across ten random seeds for both models, with per-seed results provided in Supplementary Information Table 4. The results are unambiguous: during the crash period, the model trained on data with temporal overlap predicted the initial market decline significantly better than the clean model, outperforming it in eight out of ten runs with individual improvements reaching up to approximately 37%. Up to the first of May, the temporal overlap model achieved an average MAE of about 190 compared to about 217 for the non-leakage model. For the remaining part, both models converged to very similar average performance (108.3 vs. 108.7), confirming that the advantage was specific to the period of shared causal influence rather than reflecting a general training benefit. This performance gap is all the more striking given that the leaked signal comprised just seven time series out of 484, or roughly 1,600 out of over 205,000 total training observations (Table 3). If a small transformer trained from scratch on fewer than 500 series can exploit temporal correlation for a measurable advantage, production-scale TSFMs trained on millions of series are likely more susceptible, not less, given their greater capacity to memorize subtle patterns and the higher probability that their pre-training corpora contain series that are temporally overlapped and correlated with any given test set.
Both the Madrid transport experiment (Rodrigo and Ortiz, 2024) and the COVID stock market experiment involve indirect leakage between series of the same frequency. Yet a further, largely undiscussed complication concerns the relationship between different frequency representations of the same underlying process. The community has not yet reached consensus on this matter: some treat different sampled frequencies as practically independent (Auer et al., 2025; Shchur et al., 2025; Das et al., 2024), while others leverage similar patterns across frequencies architecturally and consider them dependent (Liu et al., 2024b; Graf et al., 2025). We argue that frequency variants of the same underlying process should be treated as dependent, for two reasons. First, time series sampled at different frequencies from the same source retain very high correlation when mapped to common timestamps. Second, if a TSFM is capable of capturing the “bigger picture” of a time series (and this is precisely the promise of foundation models), it should recognize the same trends and seasonal patterns regardless of temporal resolution. Training on one frequency and testing on another from the same ground truth should therefore be considered a form of indirect leakage.
4 Two Requirements for Information-Leakage-Free TSFM Evaluation
@@
However, the paradigm shift behind TSFMs shatters this assumption. Unlike local models, TSFMs learn patterns globally. Consequently, strict temporal ordering within a single time series is no longer sufficient to guarantee a fair evaluation. To rigorously assess these models, we must address leakage across two distinct dimensions: through what was in the training data (the sample dimension) and through when the training data was observed (the temporal dimension) (see Figure 5).
Requirement 1: Full flexibility in pre-training data. Attempting to prevent contamination by restricting what data TSFMs may train on is both counterproductive and unenforceable. Restrictions artificially limit the potential of foundation models, and the lineage analysis presented here demonstrates that verifying the absence of train-test overlap across opaque, large-scale pre-training corpora is practically impossible. The responsibility for preventing leakage must therefore shift from model developers to benchmark designers. A robust evaluation framework should be agnostic to pre-training choices, allowing any available data to be used for training. This means that test sets must be genuinely novel: data that could not plausibly have appeared in any pre-training corpus. The benchmark, not the model, must guarantee separation.
Requirement 2: Strict global post-training test periods. Indirect temporal leakage cannot be prevented by controlling which series appear in training and test sets; it arises from correlation between any temporally overlapping series. A hard temporal barrier is therefore needed: if the latest pre-training timestamp across all evaluated models is $t$, then every observation in the test set must come from $t+1$ or later. This barrier must be determined globally, based on the actual training cutoff of the latest model in the comparison, rather than imposed as a fixed date. Fixing a universal cutoff would require all researchers to align their training protocols to the same date and would prevent models from learning current patterns, both of which are undesirable. A flexible, model-aware temporal barrier $t_{now}$ avoids these drawbacks while ensuring that no test observation $t_{now}+1$ could have been influenced by any training observation through shared causal drivers.
Together, these two requirements form a coherent basement for new frameworks. Requirement 1 addresses the sample dimension of leakage by making test data provably novel. Requirement 2 addresses the temporal dimension by ensuring strict temporal precedence of all training over all test data. Neither requirement alone is sufficient; both are necessary. Realizing both demands not incremental adjustments to existing evaluation practices, but a structural rethinking of how TSFM benchmarks are designed and maintained.
@@ 5 Conclusion
We call on the community to develop new benchmark methodologies grounded in both requirements proposed here: full flexibility in pre-training data and strict global post-training test periods. Possible directions include constantly newly sourced test data, synthetic test data, or fast-paced competitions on unreleased and unbiased test data.
While these two requirements define the conditions for trustworthy evaluation, several critical questions remain open. To date, the effect size of test-set contamination, when a test dataset is present in the pre-training corpus, has only been shown to be significant in isolated cases. Likewise, the magnitude of temporal leakage arising from correlated and overlapping time series remains largely unknown in real TSFM settings and is likely to vary across domains, types of global events, and other contextual factors. Carefully designed TSFM training regimes will be required to rigorously quantify these effects.
@@ 6 Data Availability
The resulting table of the TSFM data-lineage analysis is supplied in the Github Repository under this link. For the indirect information leakage experiment in Section 3.2, the used data is available through the source code as well via Github: https://github.com/DAG-UPB/rethinking-tsfm-evaluation
@@ 7 Code Availability
The code to reproduce the indirect information leakage experiment in Section B is provided via Github: https://github.com/DAG-UPB/rethinking-tsfm-evaluation
Appendix A Leakage Investigations
A.1 Cases of Overlaps in Train-Test Samples
This section summarizes individual cases where train-test sample overlaps unintentionally led to information leakage in recent benchmarking studies. Through our lineage analysis, we could easily compare the used benchmark datasets in the study against the TSFM training datasets (for full lineage table see Section 6).
@@ A.1 Cases of Overlaps in Train-Test Samples
In one case, the benchmark creators, due to the TSFMs’ incomprehensible datasets, unintentionally included three evaluation datasets as test sets that had already been used for the pretraining of TimesFM, UniTS, and TTM (Li et al., 2024b). Based on our analysis of the benchmark results, this lead to an 47% - 184% lower mean squared error (MSE) rate compared to best models not pre-trained on the leaking datasets. The advantage of the best TSFM on non-leaked datasets is only between 0.3% and 14%. This performance benefit through information leakage appears to align with findings from the Moirai leakage example (see Section A.2). In another example a peer-reviewed paper benchmarked TSFM models on the Electricity dataset, which has been used for pretraining of every of tested TSFMs Saravanan et al. (2024). Also in the energy domain, a paper benchmarked TSFMs on the Spanish dataset (among others), which is also included in the pretraining data of the tested TSFMs. Chronos achieved a sMAPE of 4.854 where the best non-TSFM model TiDE achieved a sMAPE of 8.102 Montet et al. (2025). Taken together, these illustrative case highlight the increasing difficulties for finding suitable benchmarking data for TSFMs.
A.2 Intended Information-Leakage Experiment for Moirai Model
@@
Table 2: Information leakage in TSFM comparing Moirai with (Moirai Leakage) and without (Moirai) information leakage during training phase for different model sizes. Metric is MAPE. Based on experiments by Aksu et al. (2024).
@@
Aksu et al. (2024) performed an empirical analysis using the Moirai TSFM to demonstrate the effects of information leakage between training and test sets. The authors prepared a new pre-training and held-out evaluation dataset without any overlaps and trained the Moirai architecture on it. In addition, they took an already pretrained Moirai model from a previous publication, whose pre-training dataset contained 0.1% of the newly defined held-out evaluation data, resulting in a deliberately small amount of information leakage. The leaked data includes nine datasets and three different forecast horizons (short, medium, and long), though not all combinations of datasets and horizons are present.
Figure 6 summarizes the MAPE of both models. On short horizons, Moirai Leakage has on average an 8 percentage points lower MAPE score across all model sizes. At medium horizons, Moirai Leakage exhibits an average accuracy advantage of 15 percentage points. And on long horizons, Moirai Leakage achieves an average MAPE score that is 29 percentage points lower than the leakage-free Moirai.
Evaluation across model sizes reveals that larger models benefit more substantially from data leakage. The small model (S) shows a modest average improvement of 4 percentage points, while the base model (B) demonstrates a more pronounced advantage of 13 percentage points. Most notably, the large model (L) exhibits the strongest leakage effect with an average MAPE reduction of 34 percentage points across all forecasting horizons, indicating that model capacity amplifies the impact of training-test data contamination.
These results vividly demonstrate the significant impact that train-test sample overlap can have on the performance of TSFMs and show that larger models are especially prone to this type of leakage.
@@
A.3 Potential Overlaps of Correlated Train and Test Series in Dataset Collections
It may be theoretically possible that existing model evaluation results have already been influenced by overlaps of correlated train and test series. We analyse the overlapping time frames in the Monash Dataset  [^9] https://huggingface.co/datasets/Monash-University/monash_tsf as can be seen in Figure 7. Test datasets shouldn’t lie in the same time period as other training datasets. As an example, the proposed Australian Electricity Demand test time frame (April 2015) lays in the same period as the training period of the "Weather (Australia)" training set (1900-2021). The Moirai and Moirai-MoE models use both datasets for train/test forecasts, meaning if there are any global patterns in the weather data which could influence the Australian Electricity Demand data, this would be a kind of information leakage, at least on theoretical basis. Similarly, the Oikolab Weather dataset ranges from January 2010 to May 2021, which is weather data located in Melbourne. The test set of the dataset Pedestrian Count (also Melbourne) is located at the end of April 2020. This test period lays within the training data of the Oikolab Weather, which could be a greater potential global pattern information leakage than in previous example. Both datasets are included in the datasets of Moirai, Time-Moe, Sundial and Toto.